# Using Causality for Enhanced Prediction of Web Traffic Time Series

### v1

@@ 4. Methodology
Figure 7. An overview of the proposed method, which consists of two components: the CCMPlus module and the Backbone Time Series Model. The CCMPlus module identifies causal relationships among web services, generating a CCMPlus representation that enhances the predictive performance of the Backbone Time Series Model. Per iteration refers to the entire method’s operations on a single batch of data.
Algorithm 1 Algorithmic Procedure for the CCMPlus Module
[…]
2:Output $h_{ccm}^{iter}$ and $\mathbf{M}_{cc}^{iter}$
3:Step 1: Compute $\mathbf{E}$ from $\mathbf{\tau}$ and $\tau_{w}$.
4:Initialize $\mathbf{E}$ as an empty list.
5:for each $\tau_{i}$ in $\tau$ do
6:     Compute $E_{i}\leftarrow\left\lfloor\dfrac{\tau_{w}}{\tau_{i}}\right\rfloor$
7:     if $E_{i}$ is even then
8:         $E_{i}\leftarrow E_{i}-1$
9:     end if
10:     Append $E_{i}$ to $\mathbf{E}$
11:end for
12:Step 2: calculate the $\mathbf{M}_{cc}^{iter}$ and $h_{ccm}^{iter}$.
13:for $iter\leftarrow 0,\ldots,num\_iterations-1$ do
[…]
15:     for $i\leftarrow 1\ldots n$ do
16:         $\tau_{i}\leftarrow\tau[i-1]$,  $E_{i}\leftarrow\mathbf{E}[i-1]$
17:         if $(L_{x}-\tau_{i}\times(E_{i}-1))<0$ then
18:              continue $\triangleright$ Skip if sequence too short
19:         end if
20:         Reverse the order of both $\mathbf{X}$ and $\hat{\mathbf{X}}$ along the $L_{x}$ dimension. Then reshape $\mathbf{X}$ to $\bar{\mathbf{X}}\in\mathbb{R}^{B\cdot N\times C_{in}\times L_{x}}$
21:         $\mathbf{X}_{conv}=\text{Conv1D}\bigl{(}\bar{\mathbf{X}};\,\text{kernel\_size}=E_{i},\text{dilation}=\tau_{i},\bigr{)}$, $\mathbf{X}_{conv}\in\mathbb{R}^{B\cdot N\times C_{out}\times L_{out}}$
[…]
23:              reshape $\mathbf{X}_{conv}$ to $\mathbf{X}_{ccm}\in\mathbb{R}^{B\times N\times L_{out}\times C_{out}}$, $\mathbf{Y}\leftarrow\hat{\mathbf{X}}[:,:,:L_{out}]$
24:              $dist\_matrix\leftarrow\text{PairwiseDistances}(\mathbf{X}_{ccm},\mathbf{X}_{ccm})$, $nearest\_indices\leftarrow\text{argsort}(dist\_matrix,\text{dim}=-1)$
25:              $nearest\_indices\leftarrow nearest\_indices[:,:,:,1:(C_{out}+2)]$
26:              $nearest\_distances\leftarrow\text{Gather}(dist\_matrix,nearest\_indices)$
27:              $u_{z}\leftarrow\exp\!\Bigl{(}-\,\dfrac{nearest\_distances}{nearest\_distances[:,:,:,0{:}1]+\epsilon}\Bigr{)}$, $w_{z}\leftarrow\dfrac{u_{z}}{\sum(u_{z})+\epsilon}$
28:              $Y\_nearest\leftarrow\text{GatherTargets}(\mathbf{Y},nearest\_indices)$, $\hat{Y}\leftarrow\sum(w_{z}\times Y\_nearest)$
29:              $corr\leftarrow\dfrac{\mathrm{Cov}(\hat{Y},\,\mathbf{Y})}{\sigma(\hat{Y})\,\sigma(\mathbf{Y})+\epsilon}$, $\mathbf{M}_{train}\leftarrow corr^{\top}$
30:              if $iter$0 then
31:                  $\mathbf{M}_{train}^{iter}\leftarrow m*\mathbf{M}^{iter-1}_{train}+(1-m)*\mathbf{M}_{train}$
32:              else
33:                  $\mathbf{M}_{train}^{iter}\leftarrow\mathbf{M}_{train}$
34:              end if
35:              $\widehat{\mathbf{M}_{train}}\leftarrow\text{Softmax}(\mathbf{M}_{train}^{iter},\text{dim}=-1)$, $\mathbf{M}_{cc}^{iter}(i)\leftarrow\text{Mean}(\widehat{\mathbf{M}_{train}},dim=0)$
36:         else
37:              $\mathbf{M}_{test}\leftarrow\text{Repeat}(\mathbf{M}_{cc},B)$, $\widehat{\mathbf{M}_{train}}\leftarrow\text{Softmax}(\mathbf{M}_{test},\text{dim}=-1)$, $\mathbf{M}_{cc}^{iter}(i)\leftarrow\mathbf{M}_{cc}$
38:         end if
39:         $conv\_out\_proj\leftarrow\text{LinearProjection}(\mathbf{X}_{conv})$, $\widehat{\mathbf{X}_{ccm}}\leftarrow\text{Reshape}(conv\_out\_proj,[B,\,N,\,C_{out}])$
40:         $h_{ccm}^{iter}(i)\leftarrow\widehat{\mathbf{M}_{train}}\times\widehat{\mathbf{X}_{ccm}}$
41:         append $h_{ccm}^{iter}(i)$ to multi_space_representations, append $\mathbf{M}_{cc}^{iter}(i)$ to multi_space_corrs
42:     end for
[…]
45:end for
Inspired by the concept of causal relationships in ecology, as illustrated in Figure 2, the population dynamics of rabbits influence the abundance of grass—for instance, a higher rabbit population leads to reduced grass availability. Similarly, we observe causal relationships in the web traffic patterns of different web services, as depicted in Figure 1. For example, real-world data from Google Trends indicates that an increase in web traffic for Netflix corresponds to a decrease in traffic for the office software Outlook.
@@
To leverage the latent causality between services for traffic prediction, we extend the ecology CCM theory by integrating into a neural module, referred to as CCMPlus (CCM+). As illustrated in Figure 7, the CCMPlus module leverages causal relationships among web services to generate the CCMPlus representation (Section 4.1). This representation is then concatenated with the representation produced by the Backbone Time Series Model (Section 4.2), to enhance the accuracy of web service traffic time series prediction. Finally, we introduce the optimization process in section 4.3.
4.1. CCMPlus Module
@@
In this section, we aim to derive a feature representation that captures causal relationships across web services from raw time series data, thereby improving the precision of traffic prediction. This process involves constructing the causal correlation matrix by estimating each time point of one web service time series using points from the shadow manifolds of another web service time series, and then generating the CCMPlus representation using the causal correlation matrix.
@@
The first step is to derive the initial time series embedding. Given the raw time series $\hat{\mathbf{X}}\in\mathbb{R}^{B\times N\times L_{x}}$. Note that for constructing the shadow manifold, the order of $\hat{\mathbf{X}}$ along the $L_{x}$ dimension is reversed. We first extract date features, e.g., minutes, hours and days, and then combine them with the raw time series $\hat{\mathbf{X}}$ to form the time series embedding $\mathbf{X}\in\mathbb{R}^{B\times N\times L_{x}\times C_{in}}$. Here, $L_{x}$ represents the input time series length, and $C_{in}$ denotes the representation dimension of the input.
@@
Next, we employ 1D convolution to construct shadow manifolds for each web service traffic time series. Traditional CCM algorithm relies expert to set the value of $\tau$, which might introduce human bias. To alleviate this issue, we extend it to a multi-manifold space, so as to learn the causality from diverse shadow manifold spaces. Specifically, we initialize two vectors of $\mathbf{\tau}=[\tau_{1},\ldots,\tau_{i},\ldots,\tau_{n}]$ and the corresponding $\mathbf{E}=[E_{1},\ldots,E_{i},\ldots,E_{n}]$. For simplicity, we use the $i$-th shadow manifold, defined by the pair $(\tau_{i},E_{i})$, as a representative example to explain the subsequent procedures. The steps for other shadow manifolds differ only in the index $i$. The initialized time lag $\tau_{i}$ and the corresponding embedding dimension $E_{i}$ are generated as follows:
@@
$E_{i}=\begin{cases}\left\lfloor\frac{\tau_{w}}{\tau_{i}}\right\rfloor-1,&\text{if }\left\lfloor\frac{\tau_{w}}{\tau_{i}}\right\rfloor\bmod 2=0,\\[5.0pt] \left\lfloor\frac{\tau_{w}}{\tau_{i}}\right\rfloor,&\text{if }\left\lfloor\frac{\tau_{w}}{\tau_{i}}\right\rfloor\bmod 2=1,\end{cases}$
@@
where $\tau_{w}$ is set to 100 based on prior empirical observations for shadow manifolds as established in the literature [2]. The convolution output can be derived as follows:
@@
$\mathbf{X}_{conv}=\text{Conv1D}\bigl{(}\bar{\mathbf{X}};\,\text{kernel\_size}=E_{i},\text{dilation}=\tau_{i}\bigr{)},$
where $\bar{\mathbf{X}}\in\mathbb{R}^{B\cdot N\times C_{in}\times L_{x}}$ is reshaped from $\mathbf{X}$, the channels of input and output of Conv1D are ${C_{in}}$ and ${C_{out}}$, respectively. The trajectory length of the $i$-th shadow manifold is: $L_{out}=L_{x}-\tau_{i}(E_{i}-1)$, which aligns with CCM theory Step 1 in Section 3.2.3.
@@
After obtaining the convolution output, we can derive the shadow manifold embedding $\mathbf{X}_{ccm}\in\mathbb{R}^{B\times N\times L_{out}\times C_{out}}$ reshaped from $\mathbf{X}_{conv}$, where, for each time point in $L_{out}$, the corresponding point in the shadow manifold is represented by coordinates in the $C_{out}$ dimension. To ensure consistency with the trajectory length $L_{out}$ of the shadow manifolds across different time series, we define the prediction target as $\mathbf{Y}=\hat{\mathbf{X}}[:,:,:L_{out}]$, where $\mathbf{Y}\in\mathbb{R}^{B\times N\times L_{out}}$.
4.1.1. Training Mode
Based on $\mathbf{Y}$ and $\mathbf{X}_{ccm}$, we use the points in the shadow manifold of one web service time series to predict the values of another web service time series. The predictions are then compared with the ground truth values of the target web service time series to compute the correlation coefficient for each pair. As a result, we can obtain the causal correlation matrix: $\mathbf{M}_{train}\in\mathbb{R}^{B\times N\times N}$, where each element represents the causal relationship between a pair of web services. This procedure corresponds to Steps 2–4 in Section 3.2.3.
The causal correlation matrix $\mathbf{M}_{train}$ can be efficiently computed using matrix operations on a GPU, as outlined in Algorithm 1 line 18-25. To illustrate, we use a specific element of $\mathbf{M}_{train}$ indexed as $(l,m,n)$ to demonstrate the process for quantifying causal relationships. The value $\mathbf{M}_{train}[l,m,n]$ quantifies the causal effect of the service time series $\mathbf{Y}[l,n,:]$ (denoted with $\mathbf{y}(k)$) on $\mathbf{Y}[l,m,:]$ (denoted with $x(k)$).
$P=\mathbf{X}_{ccm}[l,m,:,:]$, with $P\in\mathbb{R}^{L_{out}\times C_{out}}$, represents the points on the shadow manifold $M_{x}$, where the coordinates of the $k$-th point are given by $\underline{x}(k)=P[k,:]$. The estimation of web service time series $y(k)$ using the shadow manifold $M_{x}$ of $x(k)$ is denoted by $\hat{y}(k)\mid M_{x}$.
@@ 4.1.1. Training Mode
First, we begin by locating the contemporaneous point on $M_{x}$, $\underline{x}(k)$, and find its $C_{out}+1$ nearest neighbors. Next, denote the time indices (from closest to farthest) of the nearest neighbors of $\underline{x}(k)$ by $t_{1},...,t_{C_{out}+1}$. These time indices corresponding to nearest neighbors to $\underline{x}(k)$ on $M_{x}$ are used to identify points in time series $y$ to estimate $y(k)$ from a locally weighted mean of the $y(t_{i})$ values:
$\hat{y}(k)\mid M_{x}=\sum_{i=1}^{C_{out}+1}w_{i}y(t_{i}),$
[…]
$w_{i}=\frac{u_{i}}{\sum_{j=1}^{C_{out}+1}u_{j}},$
where
$u_{i}=\exp\left\{-\frac{d[\underline{x}(k),\underline{x}(t_{i})]}{d[\underline{x}(k),\underline{x}(t_{1})]}\right\}.$
@@
Here, $d[\underline{x}(k),\underline{x}(t_{i})]$ denotes the Euclidean distance between the two points on the shadow manifold $M_{x}$.
Finally, we calculate the correlation coefficient $\mathbf{M}_{train}[l,m,n]$ between $\hat{y}(k)$ and $y(k)$, where $k$ ranges from 00 to $L_{out}$:
$\mathbf{M}_{train}[l,m,n]=\frac{\sum_{t=0}^{L_{out}}\left(y(k)-\overline{y}(k)\right)\left(\hat{y}(k)-\overline{\hat{y}}(k)\right)}{\sqrt{\sum_{t=0}^{L_{out}}\left(y(k)-\overline{y}(k)\right)^{2}\sum_{t=0}^{L_{out}}\left(\hat{y}(k)-\overline{\hat{y}}(k)\right)^{2}}}.$
@@
To ensure consistent and stable representations, the $\mathbf{M}_{train}$ matrix is updated using a momentum-based approach [13] with the causal correlation matrix from the previous iteration, denoted as $\mathbf{M}^{iter-1}_{cc}\in\mathbb{R}^{N\times N}$, followed by a softmax operation for normalization:
(1) $\mathbf{M}_{train}^{iter}$ $=(1-m)\cdot\mathbf{M}_{train},+m\cdot\text{Repeat}(\mathbf{M}^{iter-1}_{cc},\text{dim}=0),$
${\mathbf{M}_{cc}^{iter}(i)}$ $=\text{Mean}(\text{Softmax}(\mathbf{M}_{train}^{iter},\text{dim}=-1),\text{dim}=0),$
@@
where $iter$ denotes the current iteration number, Repeat expands $\mathbf{M}^{iter-1}_{cc}$ into $\mathbb{R}^{B\times N\times N}$, $m$ is the momentum value in the range $(0,1)$, ${\mathbf{M}_{cc}^{iter}(i)}\in\mathbb{R}^{N\times N}$ corresponds to the $i$-th shadow manifold of each web service time series, effectively capturing the causal relationships among the $N$ web services.
The next step is to compute the feature representation $h_{ccm}^{iter}(i)$ incorporating the causal relationships among web services for the $i$-th shadow manifold. We first transpose the last two dimension of the manifold embedding $\mathbf{X}_{ccm}\in\mathbb{R}^{B\times N\times L_{out}\times C_{out}}$. A linear layer is then applied to reduce the dimension $L_{out}$ to $1$. This process produces an intermediate value, denoted as $\widehat{\mathbf{X}_{ccm}}\in\mathbb{R}^{B\times N\times C_{out}}$:
(2) $\widehat{\mathbf{X}_{ccm}}=\text{LinearLayer}\big{(}\text{Transpose}(\mathbf{X}_{ccm})\big{)}.$
Then,
$\widehat{\mathbf{M}_{train}}=\mathrm{Softmax}(\mathbf{M}_{train}^{iter},\,\text{dim}=-1),$
(3) $h_{ccm}^{iter}(i)=\widehat{\mathbf{M}_{train}}\cdot\widehat{\mathbf{X}_{ccm}},$
@@
where $h_{ccm}^{iter}(i)\in\mathbb{R}^{B\times N\times C_{out}}$ is the feature representation derived from the $i$-th shadow manifold. During each training iteration, the CCMPlus representation $h_{ccm}^{iter}\in\mathbb{R}^{B\times N\times C_{out}}$ and causal correlation matrix $\mathbf{M}_{cc}^{iter}\in\mathbb{R}^{N\times N}$ are computed by averaging the representations and matrices from all shadow manifolds:
(4) $h_{ccm}^{iter}$ $=\text{Mean}(\sum_{i=1}^{n}h_{ccm}^{iter}(i)),$
$\mathbf{M}_{cc}^{iter}$ $=\text{Mean}(\sum_{i=1}^{n}\mathbf{M}_{cc}^{iter}(i)),$
@@
where ${\mathbf{M}_{cc}^{iter}(i)}\in\mathbb{R}^{N\times N}$ and ${h_{ccm}^{iter}(i)}\in\mathbb{R}^{B\times N\times C_{out}}$ are the causal correlation matrix and feature representation for the $i$-th manifold space, respectively.
4.1.2. Testing Mode
Given $\mathbf{M}_{cc}\in\mathbb{R}^{N\times N}$ from the output of the training stage, it is expanded along the batch dimension to obtain $\mathbf{M}_{test}\in\mathbb{R}^{B\times N\times N}$. A softmax normalization is then applied along the last dimension to produce $\widehat{\mathbf{M}_{train}}\in\mathbb{R}^{B\times N\times N}$:
$\mathbf{M}_{test}$ $=\text{Repeat}(\mathbf{M}_{cc},\text{dim}=0),$
$\widehat{\mathbf{M}_{train}}$ $=\text{Softmax}(\mathbf{M}_{test},\text{dim}=-1).$
Based on $\mathbf{X}_{ccm}\in\mathbb{R}^{B\times N\times L_{out}\times C_{out}}$ generated by the trained model, $\widehat{\mathbf{X}_{ccm}}\in\mathbb{R}^{B\times N\times C_{out}}$ is computed using Equation 2.
Subsequently, the $h_{ccm}^{iter}(i)$ for the $i$-th shadow manifold of the web service time series is obtained using Equation 3. Since each testing iteration involves multiple shadow manifolds, the $h_{ccm}^{iter}$ is calculated as the average of the representations across all shadow manifolds, as defined in Equation 4. While ${\mathbf{M}_{cc}^{iter}}$ remains constant and is set to $\mathbf{M}_{cc}$, which is output by the training stage.
4.2. Backbone Time Series Model

### v2

@@ 4 Methodology
Figure 4: An overview of the proposed method, which consists of the CCMPlus module and the Backbone Time Series Model. The CCMPlus module identifies latent causal relationships among web services, generating a CCMPlus representation that enhances the predictive performance of the Backbone Time Series Model.
@@
Inspired by the concept of causal relationships in ecology, for example, the population dynamics of rabbits influence the abundance of grass, we analogically observe causal relationships in the web traffic patterns of different web services. As depicted in Figure 1, real-world data from Google Trends indicates that an increase in web traffic for Netflix corresponds to a decrease in traffic for the office software Outlook. To leverage the latent causality between services for traffic prediction, we extend the CCM theory by integrating it into a neural network module, referred to as CCMPlus (CCM+). It leverages causal relationships among web services to generate the CCMPlus representation (Section 4.1). This representation is used to enhance the representation produced by the Backbone Time Series Model (Section 4.2), boosting the accuracy of web service traffic time series prediction. Finally, we introduce the optimization process in Section 4.3.
4.1 CCMPlus Module
@@
In this section, we aim to derive a feature representation that captures causal relationships across web services from raw time series data, thereby improving the precision of traffic prediction. This process follows the original CCM algorithm, and constructs the causal correlation matrix by estimating each time point of one web service time series using points from the shadow manifolds of another web service time series. After that, the casualty matrix is used to generate the CCMPlus representation.
① Multi-Manifold Embedding
@@
The first step is to derive the initial time series embedding. Given the raw time series $\hat{\mathbf{X}}\in\mathbb{R}^{N\times L}$ of length $L$ for $N$ web service traffic , we first extract date features, such as minutes, hours, and days, convert them into embeddings, and then combine these embeddings with the raw time series $\hat{\mathbf{X}}$ to form the $C$-dimensional time series embedding $\mathbf{X}\in\mathbb{R}^{N\times L\times C}$. Next, we follow the original CCM algorithm to transform the time series into shadow manifold space. Given that the time index of the time-evolving variable $X(t)$ ranges from $1$ to $L$, the shadow manifold $M_{x}$ is constructed by forming lagged coordinate vectors:
$\underline{x}(t)=[X(t),X(t-\tau),X(t-2\tau),\dots,X(t-(E-1)\tau)],$ (2)
where $t\in[1+(E-1)\tau,L]$, $\tau$ represents the time lag, and $E$ denotes the embedding dimension.
@@
The traditional CCM algorithm relies on experts to set the value of $\tau$, which might introduce human bias. However, since different web time series have different characteristics, determining $\tau$ based on expert knowledge is subjective, labor-intensive and challenging. To alleviate this issue, we extend it to a multi-manifold space, so as to learn the causality from diverse shadow manifold spaces. Specifically, we initialize two vectors of $\mathbf{\tau}=[\tau_{1},\ldots,\tau_{i},\ldots,\tau_{n}]$ and the corresponding $\mathbf{E}=[E_{1},\ldots,E_{i},\ldots,E_{n}]$. For simplicity, we use the $i$-th shadow manifold, defined by the pair $(\tau_{i},E_{i})$, as an example to explain the CCMPlus procedures. The initialized time lag $\tau_{i}$ and the corresponding embedding dimension $E_{i}$ are generated as follows:
@@
$E_{i}=\begin{cases}\left\lfloor\frac{\tau_{w}}{\tau_{i}}\right\rfloor-1,&\text{if }\left\lfloor\frac{\tau_{w}}{\tau_{i}}\right\rfloor\bmod 2=0,\\[4.62497pt] \left\lfloor\frac{\tau_{w}}{\tau_{i}}\right\rfloor,&\text{if }\left\lfloor\frac{\tau_{w}}{\tau_{i}}\right\rfloor\bmod 2=1,\end{cases}$
@@
where $\tau_{w}$ is set to $100$ based on prior empirical observations for shadow manifolds as established in the literature [1]. Given the pair $(\tau_{i},E_{i})$, the corresponding calculation process of manifold embedding in Eq. 2 can be viewed as a convolution process, and can be effectively computed with the convolution neural network. Therefore, the convolution output can be derived as follows:
@@
$\mathbf{X}_{conv}=\text{Conv1D}\bigl{(}{\mathbf{X}};\,\text{kernel size}=E_{i},\text{dilation}=\tau_{i}\bigr{)},$
where $\mathbf{X}$} is reversed along the $L$ dimension, the channels of input and output of the convolution network are $C$ and $D$, respectively. The trajectory length of the $i$-th shadow manifold can be derived as $\bar{L}=L-\tau_{i}(E_{i}-1)$.
@@
After that, we can derive the shadow manifold embedding $\mathbf{X}_{ccm}\in\mathbb{R}^{N\times\bar{L}\times D}$ reshaped from $\mathbf{X}_{conv}$, where, for each time point in $\bar{L}$, the corresponding point in the shadow manifold is represented by coordinates in $D$ dimensions. Specifically, $P=\mathbf{X}_{ccm}[m,:,:]$, with $P\in\mathbb{R}^{\bar{L}\times D}$, represents the points on the shadow manifold $M_{x}$, where the coordinates of the $k$-th point are given by $\underline{x}(k)=P[k,:]$. To ensure consistency with the trajectory length $\bar{L}$ of the shadow manifolds across different time series, we define the prediction target as $\mathbf{Y}=\hat{\mathbf{X}}[:,-\bar{L}:]$, where $\mathbf{Y}\in\mathbb{R}^{N\times\bar{L}}$. $\mathbf{Y}$} is reversed along the $\bar{L}$ dimension, the $n{\text{-th}}$ target time series is $y(k)=\mathbf{Y}[n,:]$.
② Estimate $y(k)$ within Multi-Manifold Space
We use the points in the shadow manifold $M_{x}$ of one web service time series to predict the values $y(k)$ of another web service time series, denoted by $\hat{y}(k)\mid M_{x}$. This process can validate the existence of causality between the $n{\text{-th}}$ and $m{\text{-th}}$ time series of $\hat{\mathbf{X}}$.
@@
We begin by locating the contemporaneous point $\underline{x}(k)$ in $M_{x}$, and find its $D+1$ nearest neighbors, denoting their time indices (from closest to farthest) by $t_{i_{1}},...,t_{i_{D+1}}$. Note that $D+1$ is the minimum number of points needed for a bounding simplex in an $D$-dimensional space. These neighbors are used to identify points in time series $y$ to estimate $y(k)$ from a locally weighted mean of the $y(t_{i_{s}})$ values:
$\hat{y}(k)\mid M_{x}=\sum_{s=1}^{D+1}w_{i_{s}}y(t_{i_{s}}),$
where $w_{i_{s}}$ represents the weight based on the distance between $\underline{x}(k)$ and its $s{\text{-th}}$ nearest neighbor. The weights $w_{i_{s}}$ are determined by:
$w_{i_{s}}$ $=\frac{u_{i_{s}}}{\sum_{j=1}^{D+1}u_{i_{j}}},$
$u_{i_{s}}$ $=\exp\left\{-\frac{d[\underline{x}(k),\underline{x}(t_{i_{s}})]}{d[\underline{x}(k),\underline{x}(t_{i_{1}})]}\right\},$
@@
where $d[\underline{x}(k),\underline{x}(t_{i_{s}})]$ denotes the Euclidean distance between the two points in the shadow manifold $M_{x}$.
③ Momentum-Updated Correlation Coefficient Matrix
The predictions are then compared with the ground truth values of the target web service time series to compute the correlation coefficient for each pair. As a result, we can obtain the causal correlation matrix: $\mathbf{M}\in\mathbb{R}^{N\times N}$, where each element represents the causal relationship between a pair of web services.
To illustrate, we use a specific element of $\mathbf{M}$ indexed as $(m,n)$ to demonstrate the process for quantifying causal relationships. $\mathbf{M}[m,n]$ quantifies the causal effect of the $n{\text{-th}}$ time series of $\hat{\mathbf{X}}$ on the $m{\text{-th}}$ time series of $\hat{\mathbf{X}}$ and can be calculated as follows:
$\mathbf{M}[m,n]=\frac{\sum_{k=0}^{\bar{L}-1}\left(y(k)-\overline{y}(k)\right)\left(\hat{y}(k)-\overline{\hat{y}}(k)\right)}{\sqrt{\sum_{k=0}^{\bar{L}-1}\left(y(k)-\overline{y}(k)\right)^{2}\sum_{k=0}^{\bar{L}-1}\left(\hat{y}(k)-\overline{\hat{y}}(k)\right)^{2}}}.$
@@
To ensure consistent and stable representations, the $\mathbf{M}$ matrix is updated using a momentum-based approach [10] with the causal correlation matrix $\widetilde{\mathbf{M}}^{z-1}\in\mathbb{R}^{N\times N}$ from the previous iteration zz-1, followed by a softmax operation for normalization:
$\mathbf{M}^{z}$ $=(1-\lambda)\cdot\mathbf{M}+\lambda\cdot\widetilde{\mathbf{M}}^{z-1},$ (3)
$\widehat{\mathbf{M}}^{z}(i)$ $=\text{Softmax}(\mathbf{M}^{z}),$
@@
where $z$ is the current iteration number, and $\lambda$ is the momentum value in the range $(0,1)$. $\widehat{\mathbf{M}}^{z}(i)\in\mathbb{R}^{N\times N}$ captures causal relationships among the $N$ web services with the $i$-th shadow manifold of each web service time series.
④ Causality Enhanced Time Series Representation
To incorporate causality among web services into the feature representation $\mathbf{h}_{ccm}(i)$ of time series, we use the causality score to weight the feature representation $\widehat{\mathbf{X}_{ccm}}$ of the shadow manifold embedding. Specifically, we first transpose and mapping $\mathbf{X}_{ccm}$ into $\widehat{\mathbf{X}_{ccm}}\in\mathbb{R}^{N\times D}$, and then computed the causality-weighted representation:
$\mathbf{h}_{ccm}(i)$ $=\widehat{\mathbf{M}}^{z}(i)\cdot\widehat{\mathbf{X}_{ccm}}$, (4)
$\widehat{\mathbf{X}_{ccm}}$ $=\text{Transpose}(\mathbf{X}_{ccm})\cdot\mathbf{w}+b,$ (5)
@@
where $\mathbf{h}_{ccm}(i)\in\mathbb{R}^{N\times D}$ is the feature representation derived from the $i$-th shadow manifold. The CCMPlus representation $\mathbf{\widetilde{h}_{ccm}}$ and causal correlation matrix $\widetilde{\mathbf{M}}^{z}$ are computed by averaging the representations and matrices from all shadow manifolds:
$\mathbf{\widetilde{h}_{ccm}}$ $=\text{Mean}(\sum_{i=1}^{n}\mathbf{h}_{ccm}(i)),$ (6)
$\widetilde{\mathbf{M}}^{z}$ $=\text{Mean}(\sum_{i=1}^{n}\widehat{\mathbf{M}}^{z}(i)),$
@@
where $\widetilde{\mathbf{M}}^{z}\in\mathbb{R}^{N\times N}$ and $\mathbf{\widetilde{h}_{ccm}}\in\mathbb{R}^{N\times D}$ are the causal correlation matrix of the $z$-th iteration and the feature representation of the CCMPlus module respectively.
4.2 Backbone Time Series Model